A firm had total revenue of $1,500,000 last year. Total variable costs were $750,000, and fixed costs were $250,000.

The firm sells three different products, with a sales mix of 2:6:5.

Assuming the firm's cost structure and sales mix stays the same this year, how much total revenue will the firm need if it wants to earn a profit of $600,000 this year?

a.

$1,700,000

b.

$500,000

c.

Cannot be determined from this information.

d.

$1,200,000
Answer #1

Answer:

Total revenue= $1,700,000

Working notes for the above answer is as under

Sales

1,500,000

Less: Variable cost

-750,000

Contribution margin

750,000

Total Contribution margin ratio

=750,000/1,500,000

=50%

total revenue will the firm need if it wants to earn a profit of $600,000

=Fixed cost+ desired profit/ Total Contribution margin ratio

=(250,000+600,000) /50%

=850,000 /50%

=1,700,000